Diffusion Morning Report 03/07/08

Ted Berg tberg at stanford.edu
Fri Mar 7 06:15:41 PST 2008

Good Morning All,
    Looking OK this morning Tylan 1 and 2 have boat and I/o issues. LTO 
still has a leak after pull and clean. TEOS had a pumping issue. Temp 
cal on SiGe.  No other issues at this time.

  	Date: 	3/7/08 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	down 	i/o error 	  	  	 
Tylan 2 	yellow 	Boat loader issues 	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	yellow 	pump issues 	2.71/30 	  	 
Tylan 8-SiGe 	up 	temp cal-600 done working on 650 	10.63/30 	  	 
Tylan 9-Poly 	up 	  	4.38/30 	  	 
Tylan 10-Nitride 	up 	  	4.71/30 	  	 
Tylan 11-LTO 	down 	Scheduled pull and clean 	0/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	up 	  	  	  	 
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	  	  	  	 
TEL Spa 	up 	  	  	  	 
Burn Box 	Up 	  	   	  	 
Ebeam 	up 	  	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	up 	B on line 360 psi 	 changeover working well  	  	 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20080307/716a4464/attachment.html>

More information about the diffusion mailing list